# Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions

### v2

@@ 8.1 Isolated and Continuous SLR
The RWTH-PHOENIX-Weather 2014 (PHOENIX-2014) and its extended version PHOENIX-2014-T1 (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts. A semi-automated annotation protocol was proposed, leveraging existing methods for speech recognition and accelerating the process. The signers preform gestures in a constrained environment (i.e. wearing dark clothes in front of a static gray background), probably limiting the generalization of models trained to it. On the contrary, more recent efforts such as BOBSL (BOBSL-BBC) and OpenASL (OpenASL) have significantly expanded the signer base and vocabulary, with hours of British (BSL) and American Sign Language (ASL) footage collected from TV and web sources, respectively. The more unconstrained environment and the increased number of signers improve the HGR models real-world performance. The lab-recorded KETI (KETI_MDPI) comprises samples captured from two camera angles (i.e. front and side), enlarging diversity. The aim of the dataset is to interpret Korean sign language gestures that are performed in emergency situations using skeleton and RGB data, reducing the need of professionals. Another multimodal dataset collected in a laboratory is the How2Sign (How2Sign-CVPR) for ASL recognition and translation. The entire dataset was captured in the termed Green Screen studio with a single RGB camera and depth sensor, while a subset of it in the Panoptic studio using multiple cameras from various viewpoints, enabling accurate 3D pose estimation. Pose data and text transcripts are also considered in TVB-HKSL-News (TVB-HKSL-News). The samples were collected from TV news with subtitles and Optical Character Recognition was applied to extract the text sequence, while an off-the-shelf pose estimator was utilized to obtain the human keypoints. The RGB-based CSL-Daily (Back-Translation-CVPR) provides also glosses and textual transcriptions, aiding both CSLR and SLT tasks. A sign dictionary (referred to as SignDict) was also constructed to be used for other tasks, including sign spotting and segmentation. The Auslan Daily Dataset (Auslan-Daily) consists of Australian Sign Language (AUSLAN) gestures performed by humans in a television series, TV news and web videos. These samples may contain multiple persons interacting with each other, which makes the dataset suitable for in-the-wild scenarios.
@@
Other datasets for CSLR do not provide text translations. ChicagoFSWild+ (ChicagoFSWild-plus-ICCV) was collected from web videos, including ASL signs performed by 260 signers making this dataset the largest in terms of number of subjects. HKSL (SignBERT-HKSL) was captured in a laboratory environment, providing RGBD data.
@@
Datasets for IDLSR are larger compared to datasets for CSLR, as they do not require complex annotation making the data collection easier. Many of these datasets, such as CSL (CSL500), GSL iso. (ADADOGLOU), BoshporusSign22k (bosphorussign22k), AUTSL (AUTSL), KArSL (KArSL) and FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention), include pose or depth modalities, enriching multimodal learning. All of them, except for the AUTSL, were collected in a laboratory with fixed background. Datasets that do not include such modalities, can be extended to them by using off-the-shelf DL pose and depth estimators. The ASL Dataset (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET) is the only one consisting of pose sequences exclusively. MS-ASL (MS-ASL) and WLASL (WLASL) are two RGB-based datasets isolated ASL recognition collected from publicly available web videos of the deaf community, including educational resources and ASL tutorials and performed by 222 and 119 subjects respectively, increasing the diversity. However, they contain relatively few samples compared to large-scale datasets such as BSL-1K (bsl-1K), CSLD (tim-slr_csld_dataseta) and NCSL (2plus1-D-SLR_and_NCSL). For isolated Irish Sign Language (ISL) recognition, only laboratory datasets have been proposed (e.g. INCLUDE (INCLUDE), IISL2020 (IISL2020_MDPI) and FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention)), making necessary the collection of an in-the-wild dataset. For Vietnamese SLR, authors in (Multi-VSL-dinh2025sign) collecting a large-scale multi-view dataset, considering three different point of views, enhancing diversity. Other sign languages have also been considered including Argentinian (e.g. LSA64 (Lsa64argentiniansignlanguage)), Persian (e.g. RKS-PERSIANSIGN (persian_sl_dataset)) and Iranian (e.g. ISLR101 (ISLR101-IRANIAN)). The rarity of some sign languages like them mentioned previously poses obstacles to finding relevant annotators, exacerbating the data scarcity challenge, as it will be discussed later.
@@
TVB-HKSL-News (TVB-HKSL-News) 2024 CSLR/SLT RGB, Pose 7000 2 6515 HKSL 16.07h TV
Auslan-Daily (Auslan-Daily) 2023 CSLR/SLT RGB 25000 67 - AUSLAN - TV, Web
OpenASL (OpenASL) 2022 CSLR/SLT RGB 98417 220 - ASL 288h Web
CSL-Daily (Back-Translation-CVPR) 2021 CSLR/SLT RGB 20654 10 2000 CSL 23.27h Lab
BOBSL (BOBSL-BBC) 2021 CSLR/SLT RGB 1962 39 2281 BSL 1467h TV
How2Sign (green studio) (How2Sign-CVPR) 2021 CSLR/SLT RGBD, Pose 2529 11 - ASL 79.12h Lab
Corpus FinSL (Corpus-Finnish) 2020 CSLR/SLT RGB 343 21 - FinSL 14.37h -
KETI (KETI_MDPI) 2019 CSLR/SLT RGB, Pose 14672 14 524 KSL 27.99h Lab
GSL SD/SI (ADADOGLOU) 2019 CSLR/SLT RGBD 10295 7 310 GSL 9.59h Lab
PHOENIX-2014-T1 (NEURAL-SIGN-LANGUAGE-TRANSLATION-CVPR-CAMGOZ) 2018 CSLR/SLT RGB 8257 9 1066 DGS - TV
PHOENIX-2014 (PHOENIX-2014) 2014 CSLR/SLT RGB - 9 1558 DGS 10.73h TV
HKSL (SignBERT-HKSL) 2021 CSLR RGBD 2400 6 55 HKSL - Lab
ChicagoFSWild+ (ChicagoFSWild-plus-ICCV) 2019 CSLR RGB 55232 260 26 ASL - Web
Multi-VSL (Multi-VSL-dinh2025sign) 2025 IDSLR RGB, Pose 84764 30 1000 Vietnamese - Lab
BdSLW60 (BdSLW60-rubaiyeat2025bdslw60) 2025 IDSLR RGB 9307 18 60 Bangla - Lab
ISLR101 (ISLR101-IRANIAN) 2025 IDSLR RGB, Pose 4614 10 101 Iranian - -
ISW-1000 (ISW-1000-ren2025multi) 2025 IDSLR RGB, Pose, OF 10000 10 1000 CSL - Lab
FDMSE-ISL (FDMSE-hierarchicalwindowedgraphattention) 2024 IDSLR RGBD 40033 20 2002 ISL 36h Lab
LSA64 (Lsa64argentiniansignlanguage) 2023 IDSLR RGB 3200 10 64 LSA 1.9h Lab
Signsuisse Lexicon (SIGN-muller-etal-2023-findings) 2023 IDSLR RGB 18221 - - 3 (Mult) - Web
CSLD (tim-slr_csld_dataseta) 2022 IDSLR RGB 120000 30 400 CSL 66.8h Lab
NCSL (2plus1-D-SLR_and_NCSL) 2022 IDSLR RGB 90000 30 300 CSL 50.1h Lab
KArSL (KArSL) 2021 IDSLR RGBD, Pose 75300 3 502 ArSL - Lab
NMFs-CSL (NMFs-CSL) 2021 IDSLR RGB 32010 10 1.067 CSL 24.4h Lab
BosphorusSign22k (bosphorussign22k) 2020 IDSLR RGBD, Pose 22542 6 744 TSL 19h Lab
AUTSL (AUTSL) 2020 IDSLR RGBD, Pose 38336 43 226 TSL 21h Lab
INCLUDE (INCLUDE) 2020 IDSLR RGB 4287 7 263 ISL 3h Lab
IISL2020 (IISL2020_MDPI) 2020 IDSLR RGB 12100 16 11 ISL 7h Lab
RKS-PERSIANSIGN (persian_sl_dataset) 2020 IDSLR RGB 10000 10 100 Persian - Lab
BSL-1K (bsl-1K) 2020 IDSLR RGB 280000 40 1064 BSL 1060h TV
MSASL (MS-ASL) 2019 IDSLR RGB 25513 222 1000 ASL 24.65h Web
WLASL (WLASL) 2019 IDSLR RGB 21083 119 2000 ASL 14.11h Web
GSL isol. (ADADOGLOU) 2019 IDSLR RGBD 40785 7 310 GSL 6.44h Lab
SMILE (SMILE) 2018 IDSLR RGB - 30 100 DSGS - Lab
ASL Dataset (TAXONOMY_OF_HAND_GESTURES_2000_ASL_DATASET) 2018 IDSLR Pose 1200 20 30 ASL - Lab
CSL (CSL500) 2018 IDSLR RGBD, Pose 125000 50 500 CSL - Lab
@@ 8.2 Semaphoric HGR
SHGR datasets target gesture recognition from single frames, making them easier to process due to the lack of temporal complexity. Older datasets like HGR-1 (HGR-1-2024-PATTER-RECOGNITION), NUS-I (NUS-I-2010), NUS-II (NUS-II-2013), and OUHANDS (ouhands) contain few samples, limiting deep learning training, yet remain benchmarks for HCI due to their simplicity. LaRED (LaRED) offers 243K RGBD samples across 81 gestures, but only from 10 subjects, restricting subject-independent recognition. HaGRID (WACV_kapitanov2024hagrid) is much larger and diverse, with nearly twice as many full-resolution RGB frames from  35K subjects in varied scenes, including bounding boxes for hand detection. HaGRIDv2 (HAGRIDv2_ARXIV) expands this with 15 additional classes,  66K subjects, and over 1M samples. SHAPE (SHAPE_DATASET) has a similar number of gestures (all two-handed) but fewer samples. LRHG (LONG_RANGE_HGR_SSD) and MD-UHGRD (MD-UHGRD_SA-YOLO) target UAV-based gesture recognition, with MD-UHGRD offering more samples but fewer classes, captured from varying distances and angles to enhance viewpoint robustness. URGR (ULTRA-RANGE-UGV-Bamani) focuses on ultra-range gestures for UGVs, with 6 robot-command gestures performed by 16 signers at distances from 2 to 25 meters.
@@
IDHGR datasets contain gestures performed in isolation, with start and end points explicitly segmented. Most of the existing datasets (e.g. SKIG (SKIG), NVIDIA (aka nvGesture) (NVIDIA_GESTURE_R3DCNN_Molchanov), DHG (DHG-CVPRW) and SHREC17 (SHREC)) for IDHGR are relatively small with a few thousands samples and support multimodal settings by providing depth maps, IR frames, optical flow or skeleton joints, accompanying the RGB modality aiming to enhance the accuracy of the models. Jester-20bn (Jester-Dataset), proposed for gesture-based HCI, is the largest dataset for (non-SLU) IDHGR containing 1̃50K videos performed by 1376 subjects. The remarkable diversity and size of this dataset make it ideal for pretraining DL models. Other datasets for HCI include IPN-Hand (IPN-HAND), ZJUGesture (ZJUGesture) and DATE (DATE), comprising 4218, 9892 and 13500 samples respectively. LD-ConGR (LD-ConGR), ChaLearn IsoGD (chalearn-lap) and EgoGesture (EgoGesture_DATASET) have sufficient amount of data, constituting adequate baseline for IDHGR. EgoDriving (EgoFormer_METHOD_EgoDriving_DATASET), Briareo (Briareo_dataset) and nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov) have been proposed for human-car interaction, mainly in multimodal setting, including gestures that correspond to car functionalities (e.g. CW rotation and audio manipulation [^2] https://youtu.be/NJmk1DUyyB8). TCG (TCG) was also proposed for human-car interaction, but for recognizing traffic police gestures in autonomous driving scenarios relying only on pose data. EHWGesture (EHWGesture-amprimo2025ehwgesture) is the first dataset for recognizing clinical gestures and providing annotations for many other tasks, including gesture detection & triggering.
@@
CDHGR datasets challenge models to identify and segment gestures from continuous video streams. Many of the datasets mentioned earlier (e.g. ChaLearn (chalearn-lap), IPN-Hand (IPN-HAND), LD-ConGR (LD-ConGR), EgoGesture (EgoGesture_DATASET) and ZJUGesture (ZJUGesture)) have been obtained by annotating the temporal boundaries of gestures within continuous sequences, thus these datasets have protocols for both IDHGR and CDHGR. Therefore, we will only discuss the Police Gesture Dataset (Police-Gesture-Dataset) as the previous analysis could also be applied here. This dataset was introduced for recognizing police traffic gestures, similarly with TCG (TCG), containing a small number of samples (21 RGB videos) and 3354 gesture instances.
@@
HaGRID (WACV_kapitanov2024hagrid) 2024 HCI RGB 552992 34730 19 1920x1080 0.5-4m
HaGRIDv2 (HAGRIDv2_ARXIV) 2024 HCI RGB 1086158 65977 34 1920x1080 0.5-4m
MD-UHGRD (MD-UHGRD_SA-YOLO) 2024 HRI RGB 20000 - 19 - 3-10m
URGR (ULTRA-RANGE-UGV-Bamani) 2024 HRI RGB 347483 16 6 640x480 1-25m
SHAPE (SHAPE_DATASET) 2022 - RGB 34226 20 32 4128x3096 2-5m
LRHG (LONG_RANGE_HGR_SSD) 2021 HRI RGB 4320 8 10 640x480 1-7m
OUHANDS (ouhands) 2016 HCI RGBD 3150 23 10 640x480 $\leq$ 1m
HGR-1 (HGR-1-2024-PATTER-RECOGNITION) 2014 - RGB 899 12 27 174x131-640x480 $\leq$ 1m
LaRED (LaRED) 2014 - RGBD 243000 10 81 640x480 -
NUS-II (NUS-II-2013) 2013 - RGB 2000 40 10 160x120 $\leq$ 1m
ASL-FS (ASL-FS-2011-CVPRW) 2011 - RGBD 48K 4 24 - -
NUS-I (NUS-I-2010) 2010 - RGB 240 - 10 160x120 $\leq$ 1m
@@
EHWGesture (EHWGesture-amprimo2025ehwgesture) 2025 - RGBD 1100 25 11 3.3M
EgoDriving (EgoFormer_METHOD_EgoDriving_DATASET) 2024 HMI RGB 400 - 29 70276
DATE (DATE) 2024 HCI RGB 13500 22 27 -
ZJUGesture (isol.) (ZJUGesture) 2023 HCI RGB 9892 60 9 -
LD-ConGR (isol.) (LD-ConGR) 2022 HCI RGBD 44887 30 10 -
IPN Hand (isol.) (IPN-HAND) 2021 HCI RGB 4218 50 13 -
TCG (TCG) 2020 HMI Pose 250 5 4 839350
Jester (Jester-Dataset) 2019 HCI RGB 148092 1376 27 5331312
Briareo (Briareo_dataset) 2019 HMI RGBD, Pose, IR 1440 40 12 -
EgoGesture (isol.) (EgoGesture_DATASET) 2018 HCI RGBD 24161 50 83 2953224
SHREC17-14/28 (SHREC) 2017 - Depth, Pose 2800 28 14/28 -
nvGesture (NVIDIA_GESTURE_R3DCNN_Molchanov) 2016 HMI RGBD, IR, OF 1532 20 25 -
ChaLearn IsoGD (chalearn-lap) 2016 - RGBD 47933 21 249 -
DHG-14/28 (DHG-CVPRW) 2016 - Depth, Pose 2800 20 14/28 -
SKIG (SKIG) 2013 - RGBD 1080 6 10 -
@@
ZJUGesture (cont.) (ZJUGesture) 2023 HCI RGB - 9892 60 9
LD-ConGR (cont.) (LD-ConGR) 2022 HCI RGBD 542 44887 30 10
IPN Hand (cont.) (IPN-HAND) 2021 HCI RGB 200 4218 50 13
Police Gesture Dataset (Police-Gesture-Dataset) 2020 HMI RGB 21 3354 - 8
EgoGesture (cont.) (EgoGesture_DATASET) 2018 HCI RGBD 2081 24161 50 83
ChaLearn ConGD (chalearn-lap) 2016 - RGBD 22535 47933 21 249

### v3

@@ 8.1 Isolated and Continuous SLR
The RWTH-PHOENIX-Weather 2014 (49) and its extended version PHOENIX-2014-T1 (27) are benchmark datasets offering sentence-level annotations in German Sign Language (DGS) captured from television weather forecasts. A semi-automated annotation protocol was proposed, leveraging existing methods for speech recognition and accelerating the process. The signers preform gestures in a constrained environment (i.e. wearing dark clothes in front of a static gray background), probably limiting the generalization of models trained to it. On the contrary, more recent efforts such as BOBSL (12) and OpenASL (178) have significantly expanded the signer base and vocabulary, with hours of British (BSL) and American Sign Language (ASL) footage collected from TV and web sources, respectively. The more unconstrained environment and the increased number of signers improve the HGR models real-world performance. The lab-recorded KETI (101) comprises samples captured from two camera angles (i.e. front and side), enlarging diversity. The aim of the dataset is to interpret Korean sign language gestures that are performed in emergency situations using skeleton and RGB data, reducing the need of professionals. Another multimodal dataset collected in a laboratory is the How2Sign (44) for ASL recognition and translation. The entire dataset was captured in the termed Green Screen studio with a single RGB camera and depth sensor, while a subset of it in the Panoptic studio using multiple cameras from various viewpoints, enabling accurate 3D pose estimation. Pose data and text transcripts are also considered in TVB-HKSL-News (144). The samples were collected from TV news with subtitles and Optical Character Recognition was applied to extract the text sequence, while an off-the-shelf pose estimator was utilized to obtain the human keypoints. The RGB-based CSL-Daily (230) provides also glosses and textual transcriptions, aiding both CSLR and SLT tasks. A sign dictionary (referred to as SignDict) was also constructed to be used for other tasks, including sign spotting and segmentation. The Auslan Daily Dataset (177) consists of Australian Sign Language (AUSLAN) gestures performed by humans in a television series, TV news and web videos. These samples may contain multiple persons interacting with each other, which makes the dataset suitable for in-the-wild scenarios.
@@
Other datasets for CSLR do not provide text translations. ChicagoFSWild+ (179) was collected from web videos, including ASL signs performed by 260 signers making this dataset the largest in terms of number of subjects. HKSL (233) was captured in a laboratory environment, providing RGBD data.
@@
Datasets for IDLSR are larger compared to datasets for CSLR, as they do not require complex annotation making the data collection easier. Many of these datasets, such as CSL (87), GSL iso. (2), BoshporusSign22k (150), AUTSL (183), KArSL (182) and FDMSE-ISL (152), include pose or depth modalities, enriching multimodal learning. All of them, except for the AUTSL, were collected in a laboratory with fixed background. Datasets that do not include such modalities, can be extended to them by using off-the-shelf DL pose and depth estimators. The ASL Dataset (17) is the only one consisting of pose sequences exclusively. MS-ASL (94) and WLASL (110) are two RGB-based datasets isolated ASL recognition collected from publicly available web videos of the deaf community, including educational resources and ASL tutorials and performed by 222 and 119 subjects respectively, increasing the diversity. However, they contain relatively few samples compared to large-scale datasets such as BSL-1K (11), CSLD (202) and NCSL (200). For isolated Irish Sign Language (ISL) recognition, only laboratory datasets have been proposed (e.g. INCLUDE (188), IISL2020 (103) and FDMSE-ISL (152)), making necessary the collection of an in-the-wild dataset. For Vietnamese SLR, authors in (41) collecting a large-scale multi-view dataset, considering three different point of views, enhancing diversity. Other sign languages have also been considered including Argentinian (e.g. LSA64 (167)), Persian (e.g. RKS-PERSIANSIGN (163)) and Iranian (e.g. ISLR101 (160)). The rarity of some sign languages like them mentioned previously poses obstacles to finding relevant annotators, exacerbating the data scarcity challenge, as it will be discussed later.
@@
TVB-HKSL-News (144) 2024 CSLR/SLT RGB, Pose 7000 2 6515 HKSL 16.07h TV
Auslan-Daily (177) 2023 CSLR/SLT RGB 25000 67 - AUSLAN - TV, Web
OpenASL (178) 2022 CSLR/SLT RGB 98417 220 - ASL 288h Web
CSL-Daily (230) 2021 CSLR/SLT RGB 20654 10 2000 CSL 23.27h Lab
BOBSL (12) 2021 CSLR/SLT RGB 1962 39 2281 BSL 1467h TV
How2Sign (green studio) (44) 2021 CSLR/SLT RGBD, Pose 2529 11 - ASL 79.12h Lab
Corpus FinSL (173) 2020 CSLR/SLT RGB 343 21 - FinSL 14.37h -
KETI (101) 2019 CSLR/SLT RGB, Pose 14672 14 524 KSL 27.99h Lab
GSL SD/SI (2) 2019 CSLR/SLT RGBD 10295 7 310 GSL 9.59h Lab
PHOENIX-2014-T1 (27) 2018 CSLR/SLT RGB 8257 9 1066 DGS - TV
PHOENIX-2014 (49) 2014 CSLR/SLT RGB - 9 1558 DGS 10.73h TV
HKSL (233) 2021 CSLR RGBD 2400 6 55 HKSL - Lab
ChicagoFSWild+ (179) 2019 CSLR RGB 55232 260 26 ASL - Web
Multi-VSL (41) 2025 IDSLR RGB, Pose 84764 30 1000 Vietnamese - Lab
BdSLW60 (168) 2025 IDSLR RGB 9307 18 60 Bangla - Lab
ISLR101 (160) 2025 IDSLR RGB, Pose 4614 10 101 Iranian - -
ISW-1000 (165) 2025 IDSLR RGB, Pose, OF 10000 10 1000 CSL - Lab
FDMSE-ISL (152) 2024 IDSLR RGBD 40033 20 2002 ISL 36h Lab
LSA64 (167) 2023 IDSLR RGB 3200 10 64 LSA 1.9h Lab
Signsuisse Lexicon (140) 2023 IDSLR RGB 18221 - - 3 (Mult) - Web
CSLD (202) 2022 IDSLR RGB 120000 30 400 CSL 66.8h Lab
NCSL (200) 2022 IDSLR RGB 90000 30 300 CSL 50.1h Lab
KArSL (182) 2021 IDSLR RGBD, Pose 75300 3 502 ArSL - Lab
NMFs-CSL (80) 2021 IDSLR RGB 32010 10 1.067 CSL 24.4h Lab
BosphorusSign22k (150) 2020 IDSLR RGBD, Pose 22542 6 744 TSL 19h Lab
AUTSL (183) 2020 IDSLR RGBD, Pose 38336 43 226 TSL 21h Lab
INCLUDE (188) 2020 IDSLR RGB 4287 7 263 ISL 3h Lab
IISL2020 (103) 2020 IDSLR RGB 12100 16 11 ISL 7h Lab
RKS-PERSIANSIGN (163) 2020 IDSLR RGB 10000 10 100 Persian - Lab
BSL-1K (11) 2020 IDSLR RGB 280000 40 1064 BSL 1060h TV
MSASL (94) 2019 IDSLR RGB 25513 222 1000 ASL 24.65h Web
WLASL (110) 2019 IDSLR RGB 21083 119 2000 ASL 14.11h Web
GSL isol. (2) 2019 IDSLR RGBD 40785 7 310 GSL 6.44h Lab
SMILE (45) 2018 IDSLR RGB - 30 100 DSGS - Lab
ASL Dataset (17) 2018 IDSLR Pose 1200 20 30 ASL - Lab
CSL (87) 2018 IDSLR RGBD, Pose 125000 50 500 CSL - Lab
@@ 8.2 Semaphoric HGR
SHGR datasets target gesture recognition from single frames, making them easier to process due to the lack of temporal complexity. Older datasets like HGR-1 (98), NUS-I (106), NUS-II (153), and OUHANDS (132) contain few samples, limiting deep learning training, yet remain benchmarks for HCI due to their simplicity. LaRED (74) offers 243K RGBD samples across 81 gestures, but only from 10 subjects, restricting subject-independent recognition. HaGRID (95) is much larger and diverse, with nearly twice as many full-resolution RGB frames from  35K subjects in varied scenes, including bounding boxes for hand detection. HaGRIDv2 (145) expands this with 15 additional classes,  66K subjects, and over 1M samples. SHAPE (35) has a similar number of gestures (all two-handed) but fewer samples. LRHG (232) and MD-UHGRD (213) target UAV-based gesture recognition, with MD-UHGRD offering more samples but fewer classes, captured from varying distances and angles to enhance viewpoint robustness. URGR (18) focuses on ultra-range gestures for UGVs, with 6 robot-command gestures performed by 16 signers at distances from 2 to 25 meters.
@@
IDHGR datasets contain gestures performed in isolation, with start and end points explicitly segmented. Most of the existing datasets (e.g. SKIG (121), NVIDIA (aka nvGesture) (138), DHG (39) and SHREC17 (185)) for IDHGR are relatively small with a few thousands samples and support multimodal settings by providing depth maps, IR frames, optical flow or skeleton joints, accompanying the RGB modality aiming to enhance the accuracy of the models. Jester-20bn (131), proposed for gesture-based HCI, is the largest dataset for (non-SLU) IDHGR containing 1̃50K videos performed by 1376 subjects. The remarkable diversity and size of this dataset make it ideal for pretraining DL models. Other datasets for HCI include IPN-Hand (20), ZJUGesture (212) and DATE (37), comprising 4218, 9892 and 13500 samples respectively. LD-ConGR (118), ChaLearn IsoGD (197) and EgoGesture (220) have sufficient amount of data, constituting adequate baseline for IDHGR. EgoDriving (155), Briareo (128) and nvGesture (138) have been proposed for human-car interaction, mainly in multimodal setting, including gestures that correspond to car functionalities (e.g. CW rotation and audio manipulation [^2] https://youtu.be/NJmk1DUyyB8). TCG (208) was also proposed for human-car interaction, but for recognizing traffic police gestures in autonomous driving scenarios relying only on pose data. EHWGesture (16) is the first dataset for recognizing clinical gestures and providing annotations for many other tasks, including gesture detection & triggering.
@@
CDHGR datasets challenge models to identify and segment gestures from continuous video streams. Many of the datasets mentioned earlier (e.g. ChaLearn (197), IPN-Hand (20), LD-ConGR (118), EgoGesture (220) and ZJUGesture (212)) have been obtained by annotating the temporal boundaries of gestures within continuous sequences, thus these datasets have protocols for both IDHGR and CDHGR. Therefore, we will only discuss the Police Gesture Dataset (73) as the previous analysis could also be applied here. This dataset was introduced for recognizing police traffic gestures, similarly with TCG (208), containing a small number of samples (21 RGB videos) and 3354 gesture instances.
@@
HaGRID (95) 2024 HCI RGB 552992 34730 19 1920x1080 0.5-4m
HaGRIDv2 (145) 2024 HCI RGB 1086158 65977 34 1920x1080 0.5-4m
MD-UHGRD (213) 2024 HRI RGB 20000 - 19 - 3-10m
URGR (18) 2024 HRI RGB 347483 16 6 640x480 1-25m
SHAPE (35) 2022 - RGB 34226 20 32 4128x3096 2-5m
LRHG (232) 2021 HRI RGB 4320 8 10 640x480 1-7m
OUHANDS (132) 2016 HCI RGBD 3150 23 10 640x480 $\leq$ 1m
HGR-1 (98) 2014 - RGB 899 12 27 174x131-640x480 $\leq$ 1m
LaRED (74) 2014 - RGBD 243000 10 81 640x480 -
NUS-II (153) 2013 - RGB 2000 40 10 160x120 $\leq$ 1m
ASL-FS (154) 2011 - RGBD 48K 4 24 - -
NUS-I (106) 2010 - RGB 240 - 10 160x120 $\leq$ 1m
@@
EHWGesture (16) 2025 - RGBD 1100 25 11 3.3M
EgoDriving (155) 2024 HMI RGB 400 - 29 70276
DATE (37) 2024 HCI RGB 13500 22 27 -
ZJUGesture (isol.) (212) 2023 HCI RGB 9892 60 9 -
LD-ConGR (isol.) (118) 2022 HCI RGBD 44887 30 10 -
IPN Hand (isol.) (20) 2021 HCI RGB 4218 50 13 -
TCG (208) 2020 HMI Pose 250 5 4 839350
Jester (131) 2019 HCI RGB 148092 1376 27 5331312
Briareo (128) 2019 HMI RGBD, Pose, IR 1440 40 12 -
EgoGesture (isol.) (220) 2018 HCI RGBD 24161 50 83 2953224
SHREC17-14/28 (185) 2017 - Depth, Pose 2800 28 14/28 -
nvGesture (138) 2016 HMI RGBD, IR, OF 1532 20 25 -
ChaLearn IsoGD (197) 2016 - RGBD 47933 21 249 -
DHG-14/28 (39) 2016 - Depth, Pose 2800 20 14/28 -
SKIG (121) 2013 - RGBD 1080 6 10 -
@@
ZJUGesture (cont.) (212) 2023 HCI RGB - 9892 60 9
LD-ConGR (cont.) (118) 2022 HCI RGBD 542 44887 30 10
IPN Hand (cont.) (20) 2021 HCI RGB 200 4218 50 13
Police Gesture Dataset (73) 2020 HMI RGB 21 3354 - 8
EgoGesture (cont.) (220) 2018 HCI RGBD 2081 24161 50 83
ChaLearn ConGD (197) 2016 - RGBD 22535 47933 21 249